For which value or values of k, if any,does the quadratic equation in x,
(k-8)xx +. Kx. -1. =0
have exactly one solution

Answer by Jk22(389)

There is 1 solution exactly when the discriminant is zero :

The discriminant is here the quantity : k^2-4*(-1)*(k-8) = k^2+4k-32 = (k+8)(k-4)

Hence when k=-8 or +4 there is one solution.